# AEDR: Training-Free AI-Generated Image Attribution via Autoencoder Double-Reconstruction

### v1

@@ 4.3 Calibration Mechanism
The inherent complexity of images significantly influences the reconstruction losses Mario et al. (2005); Redies et al. (2012); Yu and Winkler (2013). Images with homogeneous backgrounds and low texture complexity tend to exhibit minimal variations in double-reconstruction losses, while those with dynamic backgrounds and high-complexity textures demonstrate often lead to substantial loss fluctuations, refer to Appendix E. To mitigate this intrinsic complexity bias, we propose a homogeneity-aware loss calibration mechanism defined as:
$\mathcal{H}=\sum_{i=0}^{l-1}\sum_{j=0}^{l-1}\frac{P(i,j)}{1+|i-j|},$ (4)
where $P(i,j)$ denotes the co-occurrence probability of grayscale levels $i$ and $j$ at a specified orientation and offset. The parameter $l$ represents the total number of grayscale levels (default value of 32 to ensure computational efficiency), and $|i-j|$ indicates the absolute difference in intensity between the corresponding pixel pairs. Based on this, the calibrated attribution signal is defined as follows:
@@
The calibrated metric $t^{\prime}$ effectively mitigates the intrinsic complexity bias in attribution assessment, thereby enhancing the overall attribution accuracy. Detailed experimental results, refer to Section 5.5.
Figure 4: The framework of AEDR. Our method consists of three key modules: double-reconstruction based on autoencoder (4.2), calibration mechanism (4.3), and threshold determination via kernel density estimation (4.4). AEDR employs the ratio of double-reconstruction losses, calibrated by an image homogeneity metric, as the attribution signal. The decision threshold is determined via kernel density estimation and applied for origin attribution.
4.4 Threshold Determination via Kernel Density Estimation
@@
Experimental results indicate that the calibrated attribution signal $t^{\prime}$ does not conform to any specific probability distribution across different models (see Appendix B) and may include a small number of outliers. To mitigate this, we employ the kernel density estimation (KDE) method Kim and Scott (2012), which makes no assumptions about the underlying distribution and is robust to outliers. The attribution threshold is determined using the cumulative distribution function (CDF) estimated from KDE, as formalized by:
[…]
@@
where $N$ (set to 500 in our experiments) denotes the number of samples. Specifically, 500 calibrated attribution signals from images attributed to the target model are used to estimate the distribution. The parameter $h$ represents the bandwidth of the kernel density estimation Ruckthongsook et al. (2018), and $K$ is the kernel function, for which a Gaussian kernel is used. The variable $t_{i}^{\prime}$ denotes the calibrated attribution signal of the $i$-th image. The quantity $1-\alpha$ represents the target cumulative probability.
If the calibrated attribution signal $t^{\prime}<\tau$, the test image $x$ is classified as a belonging image; otherwise, it is considered a non-belonging image. Experimental results indicate that the distribution of attribution signals varies across different generative models. Therefore, the key parameter $\alpha$ must be determined according to the specific distributional characteristics of each model.
5 Experiments and Results
This section first outlines the experimental setup (Section 5.1). We then evaluate the effectiveness and efficiency of AEDR by comparing it with existing reconstruction-based approaches, namely RONAN Wang et al. (2023) and LatentTracer Wang et al. (2024) (Section 5.2 and Section 5.3). Furthermore, we analyze the generalization ability of AEDR across different autoencoder architectures (Section 5.4). Finally, we conduct ablation studies to validate the effectiveness of each component within AEDR (Section 5.5).
5.1 Experimental Setup
Models. We evaluate AEDR on 8 text-to-image Latent Diffusion Models, including five variants of Stable Diffusion Rombach et al. (2022): Stable Diffusion v1-5 (SD1.5), Stable Diffusion v2-base (SD2base), Stable Diffusion v2-1 (SD2.1), the higher-resolution Stable Diffusion XL-base-1.0 (SDXL), and the lastet released Stable Diffusion 3.5-large (SD3.5). We also include FLUX.1-dev (FLUX) Labs (2024), based on rectified flows. All of these models employ Variational Autoencoders (VAE) Kingma and Welling (2014), and we further consider two models built upon quantized autoencoders: Kandinsky 2.1 (KD2.1) Razzhigaev et al. (2023), which uses the Modulating Quantized Vectors (MoVQ), and VQdiffusion (VQDM) Tang et al. (2022), based on Vector Quantized Variational Autoencoder (VQ-VAE). For more details, refer to Appendix A.
Dataset. We construct a dataset consisting of 9,000 images, including 1,000 real images and 8,000 AI-generated images. The real images are randomly sampled from LAION-5B Schuhmann et al. (2022). To reduce the influence of image diversity on experimental results, we employ the CLIP Interrogator Pharmapsychotic (2023) to extract prompts from the 1,000 real images. Each generative model then produces images based on these 1,000 prompts (see Appendix A). The first 500 generated images are used for hyperparameter tuning and threshold determination, while the remaining 500 are used for performance evaluation.
Our method is implemented using Python 3.10.17 and PyTorch 2.1.0. All experiments are conducted on an Ubuntu 18.04 server equipped with 8 NVIDIA RTX A6000 GPUs.
5.2 Effectiveness
In this section, we evaluate the effectiveness of AEDR for origin attribution of AI-generated images from two perspectives: (1) distinguishing belonging images from those generated by other models; (2) distinguishing belonging images from real images.
Distinguishing Belonging Images from Images Generated by Other Models. We first evaluate the effectiveness of the proposed method in distinguishing belonging images from those generated by different models. We compare our method against existing reconstruction-based approaches, RONAN Wang et al. (2023) and LatentTracer Wang et al. (2024), to assess efficiency. All methods are tested on the same evaluation dataset in 8 models. For fair comparison, all baselines use official open source implementations.
Model $\mathcal{M}_{1}$ Model $\mathcal{M}_{2}$ RONAN Wang et al. (2023) LatentTracer Wang et al. (2024) AEDR (ours)
TP FP TN FN Acc TP FP TN FN Acc TP FP TN FN Acc
@@ 5.2 Effectiveness
The experimental results are shown in Table 1, where $\mathcal{M}_{1}$ represents the target model and $\mathcal{M}_{2}$ denotes other models. Our method achieves an average attribution accuracy of 95.1%, substantially outperforming RONAN and LatentTracer, which achieve 50.3% and 70.4%, respectively. While LatentTracer attains over 90% accuracy on models such as SD1.5, SD2base, and KD2.1, its performance degrades significantly on the remaining five models. The underlying reason is that when the target model is SD1.5, SD2base, and KD2.1, the reconstruction loss is on the order of $10^{-3}$, whereas for stronger models such as FLUX Labs (2024), the reconstruction loss decreases to approximately $10^{-5}$ (see Figure 1 and Appendix F). And LatentTracer lacks any mechanism to calibrate the intrinsic complexity of images, resulting in substantial overlap between the distributions of belonging and non-belonging images, and thus a high false positive rate (FPR). These findings highlight the robustness of AEDR, which improves attribution accuracy by 24.7% over the strongest baseline, particularly on state-of-the-art latent generative models.
Distinguishing Belonging Images from Real Images. We further assess the ability of AEDR to distinguish belonging images from real images. As shown in Table 2, AEDR achieves an average accuracy of 96.9%, significantly outperforming RONAN (52.2%) and LatentTracer (66.7%). These results confirm that AEDR offers a substantial advantage—30.2% improvement over the best baseline.
Table 2: Results for distinguishing belongings and real images.
Model RONAN Wang et al. (2023) LatentTracer Wang et al. (2024) AEDR (ours)
TP FP TN FN Acc TP FP TN FN Acc TP FP TN FN Acc
@@ 5.3 Efficiency
In this section, we evaluate the computational efficiency of our method. We test all 8 models with 1,000 images each, recording the average runtime for comparison. Due to memory constraints, we use bf16 precision for FLUX Labs (2024) and SD3.5 Esser et al. (2024), while the remaining six models are run with fp32 precision. LatentTracer Wang et al. (2024) initializes gradient optimization using encoded representations of the input images, while RONAN Wang et al. (2023) uses random initialization and thus requires more optimization steps. Hence, our comparison focuses on LatentTracer as the more efficient baseline. As shown in Table 4, AEDR achieves a remarkable improvement in runtime efficiency, requiring only 1% of the runtime of LatentTracer Wang et al. (2024). This gain is primarily attributed to AEDR’s design, which replaces iterative gradient-based optimization with two simple forward passes through the autoencoder.
Table 3: Runtime on different models.
Model LatentTracer Wang et al. (2024) AEDR (ours)
SD1.5 29.85s 0.27s
SD2.1 92.09s 0.62s
SD2base 32.89s 0.25s
SD3.5 29.85s 0.69s
SDXL 162.94s 1.25s
FLUX 30.64s 0.68s
VQDM 12.58s 0.06s
KD2.1 48.14s 0.39s
Table 4: Generalization to different types of AE.
AE Type Model RONAN Wang et al. (2023) LatentTracer Wang et al. (2024) AEDR (ours)
@@
5.4 Generalization Across Autoencoder Architectures
We evaluate the generalization capability of AEDR across three distinct autoencoder architectures: VAE Kingma and Welling (2014), VQ-VAE Tang et al. (2022); Van Den Oord et al. (2017), and MoVQ Razzhigaev et al. (2023); Zheng et al. (2022). As shown in Figure 4, AEDR achieves an attribution accuracy exceeding 96% on VAE-based models. On VQ-VAE, the accuracy reaches 90.85%, and on MoVQ, it attains 82.93%. The performance degradation is mainly due to the fact that the latent space of the quantized model is discrete. The input tensor is quantized to the nearest latent code, and this quantization error makes it difficult to achieve fine-grained reconstruction. Nonetheless, AEDR outperforms LatentTracer Wang et al. (2024) on VQ-VAE, while MoVQ remains challenging, highlighting an avenue for future work.
@@ 5.5 Ablation Studies
Impact of Reconstruction Loss Metric. We evaluate the effect of different reconstruction loss metrics on attribution accuracy using the SD2.1. Four commonly used metrics are compared: Mean Absolute Error (MAE), Mean Squared Error (MSE), Structural Similarity Index (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) Choi et al. (2024); Ricker et al. (2024). As shown in Table 5.5, the attribution accuracies are 97.20% (MAE), 99.40% (MSE), 99.10% (SSIM), and 90.70% (LPIPS), respectively. Therefore, we adopt MSE as the default reconstruction loss metric in AEDR.
Table 5: Results on different loss metrics.
@@
Table 6: Effects of homogeneity calibration.
@@
Impact of Homogeneity Calibration. This section examines the impact of incorporating image homogeneity Dohmen et al. (2024) to reduce the effect of inherent image complexity on attribution accuracy. As shown in Table 6, based on homogeneity calibration, most models achieve significant performance improvements, with gains ranging from 0.18% to 9.09%, whereas FLUX and VQDM exhibit decreases in attribution accuracy of 0.15% and 1.95%. While some models exhibit marginal performance declines, the overall benefit brought by the calibration mechanism outweighs these drawbacks.
Impact of Quantile Selection. This section investigates how different quantile selections affect attribution accuracy. Due to the significant variation in the distribution of attribution signals across models (see Appendix B), we select model-specific quantile parameters. Each model is tested with 1,000 belonging images: 500 are used for threshold estimation, and the remaining 500 for evaluation. For each model, we select the value of $\alpha$ that yields the highest average accuracy across both subsets as the final quantile. The selected $\alpha$ values are summarized in Table 7. More details in Appendix B.
Table 7: The Selection of Quantile.
Model $\alpha$ Estimation Evaluation Avg Acc Model $\alpha$ Estimation Evaluation Avg Acc
SD1.5 0.01 98.64% 98.30% 98.47% SDXL 0.002 99.10% 99.49% 99.30%
0.015 98.67% 98.48% 98.58% 0.003 99.25% 99.50% 99.38%
0.02 98.28% 98.53% 98.41% 0.004 99.19% 99.41% 99.30%
SD2.1 0.002 99.67% 99.69% 99.68% FLUX 0.015 96.24% 96.57% 96.41%
0.003 99.73% 99.66% 99.70% 0.02 96.28% 96.65% 96.47%
0.004 99.64% 99.67% 99.66% 0.025 96.10% 96.69% 96.40%
SD2base 0.004 99.07% 99.00% 99.04% VQDM 0.08 90.47% 90.61% 90.54%
0.005 99.04% 99.10% 99.07% 0.085 90.38% 90.85% 90.62%
0.006 99.14% 98.96% 99.05% 0.09 90.29% 90.81% 90.55%
SD3.5 0.03 96.01% 96.43% 96.22% KD2.1 0.045 83.35% 82.51% 82.93%
0.035 96.15% 96.70% 96.43% 0.05 83.47% 82.93% 83.20%
0.04 95.76% 96.77% 96.27% 0.055 83.41% 82.79% 83.10%
6 Limitations
Although our method achieves high attribution accuracy and computational efficiency for image generation models based on continuous autoencoders Kingma and Welling (2014), its performance degrades when applied to models utilizing quantized autoencoders Van Den Oord et al. (2017); Zheng et al. (2022). This limitation reduces the generalization ability of ADER in broader application scenarios. Another limitation of our method lies in its inability to distinguish between images generated by different models that share an identical autoencoder, such as Stable Diffusion v2-1 and Stable Diffusion v2-base. However, as image generation models continue to evolve rapidly, it is increasingly common for developers to fine-tune autoencoders during training to better suit specific model architectures.
@@ 7 Conclusion
In this paper, we address the critical challenges encountered by current reconstruction-based attribution methods when dealing with SOTA generative models. We propose AEDR, a novel AI-generated image attribution approach that integrates double autoencoder reconstruction with an effective calibration mechanism based on image homogeneity. By computing the reconstruction loss ratio over two consecutive reconstructions, AEDR efficiently mitigates biases introduced by varying image complexities, resulting in significantly improved attribution performance. Extensive experiments confirm that AEDR achieves a relative improvement of 25.5% in attribution accuracy, while simultaneously reducing inference time to only 1% of that required by the leading baseline methods.
Appendix A More Details of the Used Models and Images
In this section, we provide a detailed description of the models and specific parameters used in the experiments. We also present the image generation resolutions and the corresponding types of autoencoders used in each model. More details are shown in Table 8. We then demonstrate the process of extracting prompts from real images using the CLIP Interrogator, followed by generating corresponding images with the eight models described above, as illustrated in Figure 5. Our dataset is constructed based on the parameters and procedures described above, consisting of 1,000 real images and 8,000 AI-generated images. Half of the dataset is used for threshold determination and the other half is used for performance evaluation.
Table 8: Detailed settings of the models in the experiment.
Model Type of AE Image size Num inference steps Guidance scale
Stable Diffusion v1-5 VAE $\times 512$ 512 80 7.5
Stable Diffusion v2-1 VAE $\times 768$ 768 80 7.5
Stable Diffusion v2-base VAE $\times 512$ 512 80 7.5
Stable Diffusion 3.5-large VAE $\times 1024$ 1024 35 7.5
Stable Diffusion XL-base-1.0 VAE $\times 1024$ 1024 80 7.5
FLUX.1-dev VAE $\times 1024$ 1024 35 7.5
VQdiffusion VQ-VAE $\times 256$ 256 80 7.5
Kandinsky 2.1 MoVQ $\times 512$ 512 80 7.5
(a) Extracting prompt from real image using the CLIP Interrogator.
(b) Generate images using models and the same prompt in (a).
Figure 5: Prompt extraction from real images and AI-generated image demonstration.
Appendix B More Details of Quantile Selection
Due to space constraints, Section 5.5 only presents the selected quantiles for different models. Since different models exhibit varying distributions of the attribution signal $t^{\prime}$, as shown in Figure 6, we provide a detailed analysis in this section of how different quantile selections affect attribution accuracy. For each model, 4,500 images are used for threshold estimation and another 4,500 images for evaluation, resulting in a total of 9,000 images, including both real images and images generated by the eight models. The final threshold for each model is determined based on the quantile that achieves the highest average attribution accuracy. Detailed experimental results are presented in Table 9, with the best-performing results highlighted in bold.
Figure 6: Different models exhibit varying distributions of the origin attribution signal.
Table 9: The impact of different quantile selections on attribution accuracy.
Model Alpha Estimation Acc Evaluation Acc Avg Acc Model Alpha Estimation Acc Evaluation Acc Avg Acc
SD1.5 0.01 98.64% 98.30% 98.47% SDXL 0.001 98.85% 99.49% 99.17%
0.015 98.67% 98.48% 98.58% 0.002 99.10% 99.49% 99.30%
0.02 98.28% 98.53% 98.41% 0.003 99.25% 99.50% 99.38%
0.025 97.99% 98.32% 98.16% 0.004 99.19% 99.41% 99.30%
0.03 97.88% 98.16% 98.02% 0.005 99.24% 99.42% 99.33%
0.035 97.84% 97.70% 97.77% 0.006 99.29% 99.44% 99.37%
SD2.1 0.002 99.67% 99.69% 99.68% FLUX 0.01 96.15% 96.39% 96.27%
0.003 99.73% 99.66% 99.70% 0.015 96.24% 96.57% 96.41%
0.004 99.64% 99.67% 99.66% 0.02 96.28% 96.65% 96.47%
0.005 99.64% 99.57% 99.61% 0.025 96.10% 96.69% 96.40%
0.01 99.44% 99.50% 99.47% 0.03 96.14% 96.39% 96.27%
0.015 99.26% 99.30% 99.28% 0.035 96.12% 96.12% 96.12%
SD2base 0.003 99.09% 98.94% 99.02% VQDM 0.075 90.33% 90.70% 90.52%
0.004 99.07% 99.00% 99.04% 0.08 90.47% 90.61% 90.54%
0.005 99.04% 99.10% 99.07% 0.085 90.38% 90.85% 90.62%
0.008 99.14% 98.96% 99.05% 0.09 90.29% 90.81% 90.55%
0.009 99.20% 98.90% 99.05% 0.095 90.31% 90.89% 90.60%
0.01 99.16% 98.90% 99.03% 0.1 90.41% 90.72% 90.57%
SD3.5 0.025 95.88% 96.21% 96.05% KD2.1 0.03 82.64% 81.55% 82.10%
0.03 96.01% 96.43% 96.22% 0.035 82.88% 81.95% 82.42%
0.035 96.15% 96.70% 96.43% 0.04 83.19% 82.24% 82.72%
0.04 95.76% 96.77% 96.27% 0.045 83.35% 82.51% 82.93%
0.045 95.94% 96.75% 96.35% 0.05 83.47% 82.93% 83.20%
0.05 95.76% 96.88% 96.32% 0.055 83.41% 82.79% 83.10%
Experimental results show that different models require different quantile selections. As the $\alpha$ value increases from zero, the attribution accuracy first increases and then decreases, reaching its maximum at the optimal quantile. In practical deployment, model owners can continuously expand the number of sample images to compute a more precise and model-specific attribution threshold. In addition, the attribution strength can be adjusted according to the requirements of different application scenarios.
Appendix C Reasons and Advantages of Double-Reconstruction
In this section, we explain the rationale behind selecting the Double-Reconstruction strategy and highlight its advantages. We first present the loss variation over five consecutive reconstructions on the same image (both belonging and non-belonging images) using the Stable Diffusion XL-base-1.0 model, as shown in Figure 7 and Table 10. In this experiment, the image under test is used as the input for the first reconstruction, and subsequent reconstructions use the image from the previous reconstruction as the input. The computed loss values are based on the input and output images from each reconstruction, rather than being calculated in reference to the original image. At the same time, we also compared it with the cumulative loss to show the superiority of single loss change.
Table 10: Loss for five consecutive reconstructions of belonging and non-belonging images.
Reconstruction Times Belong Image Non-Belong Image
Single Loss Cumulative Loss Single Loss Cumulative Loss
First Reconstruction 0.00030581 0.00030581 0.00164355 0.00164355
Second Reconstruction 0.00021323 0.00067133 0.00046879 0.00200841
Third Reconstruction 0.00018301 0.00110267 0.00033344 0.00253991
Fourth Reconstruction 0.00016603 0.00160316 0.00029531 0.00322278
Fifth Reconstruction 0.00016056 0.00215988 0.00025636 0.00401155
(a) Five consecutive reconstructions of the belonging image.
(b) Five consecutive reconstructions of the non-belonging image.
Figure 7: (a) and (b) show the five consecutive reconstruction processes of the belonging image and the non-belonging image. We found that in each reconstruction process, the reconstruction effect of the details is not good, such as the corners of the eyes and the scarf in the silk pocket.
The experimental results show that in both the first and second reconstructions, belonging images are reconstructed in a similar fashion, with their reconstruction losses remaining nearly identical. By contrast, non-belonging images incur a much larger loss on the first pass; after this first reconstruction, such an image is effectively projected onto the distribution learned by the autoencoder, resulting in a marked decrease in the second reconstruction loss. Accordingly, we define the attribution signal as the ratio of the first reconstruction loss to the second, which not only inherently removes any absolute offset caused by differences in image complexity but also robustly discriminates between belonging and non-belonging images.
Appendix D Two Core Theoretical Insights of AEDR
Attribution Principle. Due to the stochastic sampling process inherent in continuous autoencoders, the latent vectors corresponding to images generated by the model $M$ typically lie in high-probability regions of the autoencoder’s internal feature distribution. Although minor reconstruction errors may occur due to sampling randomness, such errors remain within an acceptable range for the autoencoder, resulting in relatively stable reconstruction losses across a double-reconstruction process.
Attribution vs. Non-attribution Discrimination. For non-belonging images, the latent vector obtained during the initial reconstruction tends to deviate significantly from the learned distribution of autoencoder, leading to a higher reconstruction loss. However, after the first reconstruction(that is, feeding the first reconstructed output into the autoencoder a second time), the second output becomes better aligned with the internal feature distribution, causing a notable reduction in the second reconstruction loss. This results in a large difference between the two reconstruction losses, which serves as a discriminative signal.
Appendix E Two Types of Attribution Errors
(a) Non-Belonging Images with Small Loss Variation
(b) Belonging Images with Large Loss Variation
Figure 8: (a) and (b) respectively illustrate two types of attribution errors. The target model used for evaluation is Stable Diffusion v1-5. The test set consists of 500 images generated by Stable Diffusion v1-5 (Belonging) and 500 images generated by Kandinsky 2.1 (Non-Belonging).
Figure 8 presents two types of misattributed samples: false positives (non-belonging images identified as belonging) and false negatives (belonging images identified as non-belonging). We observe that images with low reconstruction loss ratio tend to have simple, homogeneous backgrounds and limited detail, while those with high ratios often contain complex textures and rich visual content. These examples illustrate a known limitation of autoencoder-based reconstructions in preserving fine-grained details Choi et al. [2024], Ricker et al. [2024], highlighting the necessity of incorporating a loss calibration mechanism.
Appendix F Loss Characteristics based on Gradient Reconstruction
With the rapid advancement of image generation models, gradient-based reconstructions have become nearly indistinguishable from the original images, resulting in extremely low reconstruction losses. Consequently, slight variations caused by inherent image complexity can lead to significant overlap in the loss distributions of attributed and non-attributed images, hindering precise attribution. The Figure 9 and Table 11 illustrate this issue using four test images and their reconstruction losses across eight different generative models.
Figure 9: 4 sample images for gradient reconstruction.
Table 11: Loss of gradient-based reconstruction.
Model reconstruction loss
Image-1 Image-2 Image-3 Image-4
SD1.5 0.00021821 0.00071571 0.00151912 0.00012126
SD2.1 0.00004312 0.00018105 0.00034466 0.00003273
SD2base 0.00017373 0.00067367 0.00143998 0.00010398
SD3.5 0.00008450 0.00011302 0.00013894 0.00014620
SDXL 0.00001690 0.00007892 0.00014162 0.00001657
FLUX 0.00009152 0.00008627 0.00008035 0.00005152
VQDM 0.00005064 0.00006152 0.00004454 0.00001033
KD2.1 0.00016612 0.00057640 0.00120155 0.00007604
As shown in the Table 11, consider Image-1 as an example: for lower-capacity models such as SD1.5 and SD2base, the reconstruction loss is on the order of $10^{-3}$; for more advanced models like FLUX Labs [2024] and SD3.5 Esser et al. [2024], it drops to the order of $10^{-5}$. As model performance improves, reconstruction loss continues to decrease, making the loss increasingly sensitive to inherent image complexity. This shift causes greater overlap between the distributions of belonging and non-belonging images.

### v2

@@ 3.3 Calibration Mechanism
The inherent visual complexity of an image can significantly affect its reconstruction loss. Images with homogeneous backgrounds and low texture variation typically exhibit minimal differences in double-reconstruction losses. In contrast, images containing complex textures or dynamic backgrounds often lead to substantial fluctuations in these losses (error examples see Technical Appendix, Section 6). To address this intrinsic bias, we introduce a homogeneity-aware loss calibration mechanism, defined as:
$\mathcal{H}=\sum_{i=0}^{\ell-1}\sum_{j=0}^{\ell-1}\frac{P(i,j)}{1+|i-j|},$ (4)
where $P(i,j)$ denotes the co-occurrence probability of grayscale levels $i$ and $j$ at a specified orientation and offset. The parameter $\ell$ represents the total number of grayscale levels (default value of 32 to ensure computational efficiency), and |ij||i-j| indicates the absolute difference in intensity between the corresponding pixel pairs. Based on this, the calibrated attribution signal is defined as follows:
@@
The calibrated metric $t^{\prime}$ effectively mitigates the intrinsic complexity bias in attribution assessment, thereby enhancing the overall attribution accuracy of AEDR.
3.4 Threshold Determination
@@
Experimental results show that the calibrated attribution signal $t^{\prime}$ does not follow a consistent or well-defined probability distribution across different models (see Technical Appendix, Section 5), and may contain a small number of outliers. To address this, we adopt kernel density estimation (KDE) (Kim and Scott 2012), a non-parametric method that makes no assumptions about the underlying distribution and is inherently robust to outliers. The attribution threshold is then derived from the cumulative distribution function (CDF) estimated via KDE, as defined by:
$\tau=\inf\Bigl\{u\;\Bigm|\;\int_{-\infty}^{u}\frac{1}{N\,h}\sum_{i=1}^{N}K\!\Bigl(\frac{y-t^{\prime}_{i}}{h}\Bigr)\,\mathrm{d}y\;\geq\;1-\alpha\Bigr\},$ (6)
@@
where $N$ (set to 500 in our experiments) denotes the number of samples. Specifically, we estimate the distribution using 500 calibrated attribution signals obtained from belonging images. The parameter $h$ represents the bandwidth of the kernel density estimation (Ruckthongsook et al. 2018), and $K$ is the kernel function, for which a Gaussian kernel is used. The variable $t_{i}^{\prime}$ denotes the calibrated attribution signal of the $i$-th image. The quantity $1-\alpha$ specifies the target cumulative probability for threshold selection. If the calibrated attribution signal $t^{\prime}<\tau$, the $x$ is classified as a belonging image; otherwise, it is considered a non-belonging image.
$\mathcal{M}_{1}$ $\mathcal{M}_{2}$ RONAN LatentTracer AEDR
TP FP FN TN Acc TP FP FN TN Acc TP FP FN TN Acc
@@
Model RONAN LatentTracer AEDR (ours)
TP FP FN TN Acc TP FP FN TN Acc TP FP FN TN Acc
@@
Table 2: Results for distinguishing belongings and real images. Real images are randomly selected from LAION-5B.
4 Experiments and Results
We begin by describing the experimental setup. Next, we assess the effectiveness and efficiency of AEDR in comparison to existing reconstruction-based passive attribution methods, specifically RONAN (Wang et al. 2023) and LatentTracer (Wang et al. 2024). We then evaluate the generalization capability of AEDR across various autoencoder architectures. Finally, we perform ablation studies to examine the contribution of each component within AEDR.
4.1 Experimental Setup
Models. We evaluate AEDR on eight text-to-image Latent Diffusion Models, including five Stable Diffusion variants: SD1.5, SD2base, SD2.1, SDXL, and the latest SD3.5. We also include FLUX.1-dev (FLUX), which uses rectified flows. All models use Variational Autoencoders (VAE), except for two based on quantized autoencoders: Kandinsky 2.1 (KD2.1) with Modulating Quantized Vectors, and VQdiffusion (VQDM) using Vector Quantized VAE.
Dataset. We create a dataset of 9,000 images, comprising 1,000 real images sampled from LAION-5B (Schuhmann et al. 2022) and 8,000 AI-generated images. To minimize image diversity impact, we use the CLIP Interrogator (Pharmapsychotic 2023) to extract prompts from the real images. These prompts are then used to generate images with all eight models. The first 500 generated images are for hyperparameter tuning and threshold setting, while the remaining 500 are for performance evaluation. Further details are in Section 4 of the Technical Appendix.
4.2 Effectiveness
Distinguishing Belonging Images from Images Generated by Other Models. We compare our method against existing reconstruction-based passive attribution approaches, RONAN (Wang et al. 2023) and LatentTracer (Wang et al. 2024), to assess efficiency. All methods are tested on the same evaluation dataset across 8 models. For fair comparison, all baselines use official open source implementations.
The experimental results are shown in Table 1, where $\mathcal{M}_{1}$ represents the target model and $\mathcal{M}_{2}$ denotes other models. Our method achieves an average attribution accuracy of 95.1%, substantially outperforming RONAN and LatentTracer, which achieve 50.3% and 70.4%, respectively. While LatentTracer attains over 90% accuracy on models such as SD1.5, SD2base, and KD2.1, its performance degrades significantly on the remaining five models. The underlying reason is that when the target model is SD1.5, SD2base, and KD2.1, the reconstruction loss is on the order of $10^{-4}$, whereas for stronger models such as FLUX, the reconstruction loss decreases to approximately $10^{-5}$ (see Figure 1 and Sections 1 to 3 of the Technical Appendix). And LatentTracer lacks any mechanism to calibrate the intrinsic complexity of images, resulting in substantial overlap between the distributions of belonging and non‑belonging images, and thus a high false positive rate (FPR). These findings highlight the robustness of AEDR, which improves attribution accuracy by 24.7% over the strongest baseline.
Distinguishing Belonging Images from Real Images. We further assess the ability of AEDR to distinguish belonging images from real images. As shown in Table 2, AEDR achieves an average accuracy of 96.9%, significantly outperforming RONAN (52.2%) and LatentTracer (66.7%). These results confirm that AEDR offers a substantial advantage—30.2% improvement over the best baseline. Notably, both baselines exhibit high FPR, whereas AEDR maintains a low FPR alongside its superior accuracy.
Model SD1.5 SD2.1 SD2base SD3.5
LatentTracer 29.85s 92.09s 32.89s 29.85s
AEDR(ours) 0.27s 0.62s 0.25s 0.69s
[…]
Model SDXL FLUX VQDM KD2.1
LatentTracer 162.94s 30.64s 12.58s 48.14s
AEDR(ours) 1.25s 0.68s 0.06s 0.39s
[…]
Table 3: Running time on different models.
@@ 4.3 Efficiency
[…]
4.4 Generalization
We evaluate the generalization capability of AEDR across three distinct autoencoder architectures: VAE, VQ-VAE, and MoVQ. Table 4 shows that AEDR achieves an attribution accuracy exceeding 96% on VAE-based models. On VQ-VAE, the accuracy reaches 90.85%, and on MoVQ, it drops to 82.93%. The performance degradation primarily stems from the discrete nature of the latent space in quantized models. Specifically, the input tensor is mapped to the nearest latent code during quantization, introducing an error that hampers precise reconstruction and, consequently, fine-grained attribution. Despite this, AEDR outperforms LatentTracer on VQ-VAE. MoVQ remains a more challenging case, pointing to a promising direction for future research.
AE Type Model RONAN LatentTracer AEDR
@@ 4.4 Generalization
Table 4: Generalization to different types of AE.
@@ 4.5 Ablation Studies
Impact of Reconstruction Loss Metric. We evaluate the effect of different reconstruction loss metrics on attribution accuracy using the Stable Diffusion v2-1. We utilized 500 belonging images and 500 randomly selected non-belonging images, including both generated and real images. Four commonly used metrics are compared: Mean Absolute Error (MAE), Mean Squared Error (MSE), Structural Similarity Index (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS). As shown in Table 5, the attribution accuracies are 97.20% (MAE), 99.40% (MSE), 99.10% (SSIM), and 90.70% (LPIPS), respectively. Therefore, we adopt MSE as the default reconstruction loss metric in AEDR.
@@
Table 5: Results on different loss metrics.
Impact of Homogeneity Calibration. This section investigates the effect of homogeneity calibration in mitigating the influence of inherent image complexity on attribution accuracy. As reported in Table 6, homogeneity calibration leads to notable performance gains for most models, with improvements ranging from 0.18% to 9.09%. However, FLUX and VQDM (Zheng et al. 2022) exhibit slight decreases of 0.15% and 1.95%. Despite minor performance drops in a few cases, the overall advantage of the calibration mechanism clearly outweighs these limitations.
@@
Table 6: Effects of homogeneity calibration.
Impact of Quantile Selection. This section investigates how different quantile selections affect attribution accuracy. Due to the significant variation in the distribution of attribution signals across models, we select model-specific quantile parameters. Each model is tested with 1,000 belonging images: 500 are used for threshold estimation, and the remaining 500 for evaluation. For each model, we select the value of $\alpha$ that yields the highest attribution accuracy on the estimation image set. These selected $\alpha$ values are summarized in Table 7; further details in Technical Appendix Section 5.
Model Alpha Estimation Evaluation
SD1.5 0.015 98.67% 98.48%
SD2.1 0.003 99.73% 99.66%
SD2base 0.005 99.20% 98.90%
SD3.5 0.035 96.15% 96.70%
SDXL 0.003 99.29% 99.44%
FLUX 0.02 96.28% 96.65%
VQDM 0.085 90.47% 90.61%
KD2.1 0.05 83.47% 82.93%
Table 7: The Selection of quantile selection.
@@ 5 Conclusion
[…]
Acknowledgements
This work was supported in part by the National Natural Science Foundation of China under Grant 62472398, U2336206, 62121002, 62402469 and 62302146, and by the Fundamental Research Funds of a State Key Laboratory under CBCT2024KF01.